Rebecca's mother is 4 times as old as Rebecca is now. In 6 years, Rebecca's mother will be 3 times as old as Rebecca will be then. What is Rebecca's age now, in years?
Answer by oberobic(2304)

M = 4*R
.
In 6 years
M+6
R+6
.
(M+6) = 3*(R+6)
.
substitute
.
(4*R +6) = 3*(R+6)
.
4R +6 = 3R + 18
.
R = 12
M = 4*12 = 48
.
check in 6 years
M+6 = 54
12+6 = 18
3*18 = 54
Correct
.
Answer: Rebecca is 12 years old.
.
Done.